Ballas Co. uses a periodic inventory system. Its records show the following for the month of May, in which 148 units were sold.

Unit Units Cost Total Cost May 1 Inventory65 15 Purchases55 24 Purchases 76 196 $9 12 13 $585 660 988 $2,233 Totals Compute the ending inventory at May 31 and cost of goods sold using the FIFO and LIFO methods. FIFO LIFO Ending inventory at May 31 Cost of goods sold
Answer #1

Answer

FIFO:
Ending inventory at May 31 624 =48*13
Cost of goods sold 1609 =2233-624
LIFO:
Ending inventory at May 31 432 =48*9
Cost of goods sold 1801

the beginning use total amt of units and sub by the units sold ie. 196 -148=48